# Multi-period Stochastic Network Design for Combined Natural Gas and Hydrogen Distribution

### v3

@@ 6 Case Study and Discussion: Transitioning Natural Gas Network to Hydrogen in the HEAVENN Region
Figure 4: Initial natural gas network
@@ 6.1 Results
Figure 5: Hydrogen networks selected at $t=T$
@@ 7 Conclusions
Several suggestions arise for future studies on the network design topic. Firstly, exploring multi-stage recourse models for this problem could offer a more nuanced approach to decisions across different periods. Second, considering heterogeneous customers might be practical, leading to a different penalty cost per customer. On such a setting, considering the implementation of risk-adjusted performance metrics such as Conditional Value-at-Risk (CVaR) could provide a more accurate reflection of decision-makers’ risk aversion, particularly for customers involving extreme losses. Third, incorporating uncertainty in pipeline construction or repurposing costs could be beneficial, as costs can vary based on geographical and geological conditions. Lastly, integrating principles of land-use planning or urban development might offer insights into how the strategic positioning of hydrogen refueling stations or households influences energy costs in the network.
Another promising direction involves adapting our progressive hedging-based matheuristic to other two-stage recourse models, due to its generability. Progressive hedging algorithms often exhibit late convergence when only a small subset of decision variables remain non-convergent for a substantial number of iterations (Watson and Woodruff, 2011). For this, our heuristic refines the progressive hedging algorithm by applying problem-specific constraints that recognize the similarities among decision variable subsets, rather than relying solely on individual variable convergence. These constraints manage cases where individual decision variables may differ, yet a consistent pattern across the scenarios may be utilized to accelerate convergence while maintaining high solution quality. Similarly, the small subset of decision variables remaining non-convergent may be addressed more efficiently using state-of-the-art optimization solvers such as Gurobi, as we demonstrated with an early convergence check in our algorithm. Future studies should explore this progressive hedging framework for two-stage recourse models, by combining early convergence checks, scenario bundling, and creating problem-specific heuristic constraints on subproblems, for accelerated, high-quality solutions.
@@ Appendix A Instance Generation
We further generate demand samples, $s_{it}^{k\omega}$, by considering that $\mathscr{K}=2$ with hydrogen and natural gas represented by indices $k=0$ and $k=1$, respectively. We introduce $pr_{t}$ as the probability of observing $s_{it^{\prime}}^{1\omega}=0$ for $t^{\prime}\in\{t+1,\dots,T\}$ for natural gas, and $s_{it^{\prime}}^{0\omega}=0$ for $t^{\prime}\in\{0,\dots,t-1\}$ for hydrogen. Essentially, $pr_{t}$ depicts the transition probability from natural gas to hydrogen at time $t^{\prime}$. For each scenario $\omega$, we select $pr_{t}$ such that $pr_{t}>>pr_{t-1}$ in order to represent the expected exponentiality of the transition from natural gas to hydrogen, where $pr_{0}=0$ and $\sum_{t\in\mathscr{T}}pr_{t}=1$. Specifically, we select $pr_{t}$ as;
$pr_{t}=\frac{r^{t}-1}{\sum_{t=0}^{T-1}(r^{t}-1)}$ (46)
where $r$ is uniformly randomized in $[1.01,1.4]$, differently per each scenario of each instance. Using the probability array $pr_{t}$, we randomly assign a transition time matrix, $a_{ik}$, indicating the period when commodity $k$ transitions on node $i$. For example, if $s_{it^{\prime}}^{k\omega}=0$ for $t^{\prime}\in\{t+1,\dots,T\}$ for natural gas, then $a_{ik}=t$, where $\mathbb{P}\{a_{ik}=t\}=pr_{t}$. For our base scenario, the net supply levels at the beginning are represented by $s_{i0}^{k}$, structured to account for two distinct seasons: winter and summer. Winter mostly has heightened demand but diminished supply, whereas summer experiences the opposite. Each period $t$ alternates between these seasons. The initial net supply levels are given in the Table 4.
@@
The generation of $s_{it}^{k\omega}$ is further detailed in Algorithm 5. In this methodology, the growth factor $G$ represents the expected demand increase ratio on each period over the planning horizon. We introduce an additional factor, $H$, specifically for hydrogen to capture the potentially elevated growth rates due to its emergence as a new energy source. Additionally, we introduce $R$ to account for the uncertainty in short-term energy demand fluctuations. We pick values $[G_{L},G_{U}]=[1.0,1.1]$, $[H_{L},H_{U}]=[1.3,1.7]$, and $R=0.2$ in our computational experiments.
@@
1:for each scenario $w\in\mathscr{W}$ do
2:     Randomize a probability array $pr_{t}$ for $t\in\mathscr{T}$.
3:     Assign transition time matrix; $a_{ik}$, according to $pr_{t}$.
4:     Randomize uniformly $G\in[G_{L},G_{U}]$ and $H\in[H_{L},H_{U}]$.
5:     For natural gas, assign $s_{it}^{k}=s_{i0}^{k}G^{t}$ for each node $i$ for $t<a_{ik}$, and 00 for $t\geq a_{ik}$ by the corresponding season.
6:     For hydrogen, assign $s_{it}^{k}=s_{i0}^{k}G^{t}H$ for each node $i$ for $t\geq a_{ik}$, and 00 for $t<a_{ik}$ by the corresponding season.
7:     For all node $i$, commodity $k$ and period $t$, select $s_{it}^{k\omega}$ in $[(1-R\frac{t}{T})s_{it}^{k},(1+R\frac{t}{T})s_{it}^{k}]$ uniformly random.

### v4

@@ 6 Case Study and Discussion: Transitioning Natural Gas Network to Hydrogen in the HEAVENN Region
Figure 4: The initial configuration of the natural gas network for the HEAVENN case study. The letters on the pipeline arcs represent their capacity: S for Small, M for Medium, and L for Large.
@@ 6.1 Results
Figure 5: Comparison of final network designs (at $t=T$) for the stochastic (top row) and deterministic (bottom row) solutions across the two supply cases. Green pipelines are for hydrogen, while orange ones are for natural gas. Arc labels (S, M, L) indicate capacity class.
@@ 7 Conclusions
Several suggestions arise for future studies on the network design topic. Firstly, a valuable extension would be to explore an adaptive multi-stage recourse model (Kayacık et al., 2025) in the context of multi-period stochastic network design, which could permit corrective investment decisions at a few critical junctures. This would offer a more nuanced approach than a pure two-stage model while still respecting the real-world constraints that make a multi-stage recourse impractical. Second, the scenario bundling strategy used in our matheuristic can be strengthened. While this study employs random assignment as a baseline, future research should explore deterministic clustering approaches—such as k-means based on demand profiles or flow patterns—to improve convergence behavior and solution quality. Third, considering heterogeneous customers might be practical, leading to a different penalty cost per customer. In such a setting, considering the implementation of risk-adjusted performance metrics such as Conditional Value-at-Risk (CVaR) could provide a more accurate reflection of decision-makers’ risk aversion, particularly for customers involving extreme losses. Moreover, incorporating uncertainty in pipeline construction or repurposing costs could be beneficial, as costs can vary based on geographical and geological conditions. Additionally, integrating principles of land-use planning or urban development may offer insights into how the strategic positioning of hydrogen refueling stations or households affects energy costs in the network.
Another promising future research direction involves adapting our progressive hedging-based matheuristic to other two-stage recourse models, due to its generality. Progressive hedging algorithms often exhibit late convergence when only a small subset of decision variables remains non-convergent for a substantial number of iterations (Watson and Woodruff, 2011). For this, our heuristic refines the progressive hedging algorithm by applying problem-specific constraints that recognize the similarities among decision variable subsets, rather than relying solely on individual variable convergence, i.e., aggregate soft fixing. These constraints manage cases where individual decision variables may differ, yet a consistent pattern across the scenarios may be utilized to accelerate convergence while maintaining high solution quality. Similarly, the small subset of decision variables remaining non-convergent may be addressed more efficiently using state-of-the-art optimization solvers such as Gurobi, as we demonstrated with an early convergence check in our algorithm. Future studies should explore this progressive hedging framework for two-stage recourse models, by combining aggregate soft fixing, early convergence checks, and scenario bundling, for accelerated, high-quality solutions.
@@ Appendix A Instance Generation
We further generate demand samples, $s_{it}^{k\omega}$, by considering that $\mathscr{K}=2$ with hydrogen and natural gas represented by indices $k=0$ and $k=1$, respectively. We introduce $pr_{t}^{\omega}$ as the probability of observing $s_{it^{\prime}}^{1\omega}=0$ for $t^{\prime}\in\{t+1,\dots,T\}$ for natural gas, and $s_{it^{\prime}}^{0\omega}=0$ for $t^{\prime}\in\{0,\dots,t-1\}$ for hydrogen. Essentially, $pr_{t}^{\omega}$ depicts the transition probability from natural gas to hydrogen at time $t^{\prime}$. For each scenario $\omega$, we select $pr_{t}^{\omega}$ such that $pr_{t}^{\omega}$ >>>> $pr_{t-1}^{\omega}$ in order to represent the expected exponentiality of the transition from natural gas to hydrogen, where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}pr_{0}^{\omega}}=0$ and $\sum_{t\in\mathscr{T}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}pr_{t}^{\omega}}=1$. Specifically, we select $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}pr_{t}^{\omega}$} as;
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}pr_{t}^{\omega}}=\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r(\omega)^{t}}-1}{\sum_{t=0}^{T-1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r(\omega)^{t}}-1)}$ (49)
where $r$ is uniformly randomized in $[1.01,1.4]$, differently per each scenario of each instance. Using the probability array $pr_{t}^{\omega}$ , we randomly assign a transition time matrix, $a_{ik}^{\omega}$, indicating the period when commodity $k$ transitions on node $i$. For example, if $s_{it^{\prime}}^{k\omega}=0$ for $t^{\prime}\in\{t+1,\dots,T\}$ for natural gas, then ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}a_{ik}^{\omega}}=t$, where $\mathbb{P}\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}a_{ik}^{\omega}}=t\}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}pr_{t}^{\omega}}$. For our base scenario, the net supply levels at the beginning are represented by $s_{i0}^{k}$, structured to account for two distinct seasons: winter and summer. Winter mostly has heightened demand but diminished supply, whereas summer experiences the opposite. Each period $t$ alternates between these seasons. The initial net supply levels are given in the Table 5.
@@
The generation of $s_{it}^{k\omega}$ is further detailed in Algorithm 5. In this methodology, the growth factor $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G^{\omega}$} represents the expected demand increase ratio on each period over the planning horizon. We introduce an additional factor, $H$, specifically for hydrogen to capture the potentially elevated growth rates due to its emergence as a new energy source. Additionally, we introduce $R$ to account for the uncertainty in short-term energy demand fluctuations. We pick values $G_{L},G_{U}]=[1.0,1.1]$, $H_{L},H_{U}]=[1.3,1.7]$, and $R=0.2$ in our computational experiments.
@@
1:for each scenario $\omega\in\Omega$ do
2:  Randomize a probability array $pr_{t}^{\omega}$ for $t\in\mathscr{T}$.
3:  Assign transition time matrix; $a_{ik}^{\omega}$, according to $pr_{t}^{\omega}$.
4:  Randomize uniformly ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G^{\omega}}\in[G_{L},G_{U}]$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H^{\omega}}\in[H_{L},H_{U}]$.
5:  For natural gas, assign $s_{it}^{k}=s_{i0}^{k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G^{\omega}_{t}}$ for each node $i$ for $t<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}a_{ik}^{\omega}}$, and 0 for $t\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}a_{ik}^{\omega}}$ by the corresponding season.
6:  For hydrogen, assign $s_{it}^{k}=s_{i0}^{k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G^{\omega}_{t}H^{\omega}}$ for each node $i$ for $t\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}a_{ik}^{\omega}}$, and 0 for $t<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}a_{ik}^{\omega}}$ by the corresponding season.
7:  For all node $i$, commodity $k$ and period $t$, select $s_{it}^{k\omega}$ in $(1-R\frac{t}{T})s_{it}^{k},(1+R\frac{t}{T})s_{it}^{k\omega}]$ uniformly random.
@@
To illustrate Algorithm 5, Figure 6 shows the scenario-averaged net supply and demand over time for representative nodes for both commodities. The plots demonstrate the emergent gradual transition from natural gas to hydrogen. The trend shows the expected decrease in natural gas and increase in hydrogen over the planning horizon, while the oscillations, particularly visible for the natural gas, correspond to the modeled seasonality between winter and summer periods.
022446688101012121414161605050100100150150200200250250300300350350400400450450Time PeriodHydrogen Supply02020404060608080100100120120140140160160180180200200220220240240260260Natural Gas Supply
(a) Examples of net supply transitions for selected supply nodes
0224466881010121214141616110-110100-10090-9080-8070-7060-6050-5040-4030-3020-2010-1001010Time PeriodHydrogen Demand50-5045-4540-4035-3530-3025-2520-2015-1510-105-50Natural Gas Demand
(b) Examples of net supply transitions for selected demand nodes
Figure 6: Illustration of the emergent gradual transition for average net supply (a) and demand (b) for $T=15$ and $W=120$. The overall trends show the shift from natural gas to hydrogen, while the oscillations reflect modeled seasonality.
@@ Appendix B Benders decomposition
Initially, we explored the use of Benders decomposition as it is common for stochastic network design problems (Crainic et al., 2021; Rahmaniani et al., 2018, 2024). Preliminary experiments, however, indicated significant limitations in terms of computational efficiency for our problem. In order to handle with these inefficiencies, Crainic et al. (2021) accelerate a partial Benders decomposition with strategies such as adding cuts, MIP gaps and time limits for a stochastic network design study, resulting in optimality gaps about $1-5\%$ for problems mostly sized at $N=10$ with no multi-period setting. Similar MIP gaps, and time limits are also implemented in the other Benders decomposition based network design literature (Rahmaniani et al., 2018, 2024), resulting in sub-optimal solutions. Because of this, as the problem complexity increases, particularly with multi-period settings, it is usual to apply heuristics rather than exact methods (Boland et al., 2017; Jiang et al., 2021; Fragkos et al., 2021). Overall, consistent with these approaches, we believe our algorithm offers a balance of solution quality and computational time for our stochastic network design setting, while providing optimality gaps on par with those achieved with Benders decomposition. Notably, as Benders decomposition, progressive-hedging is an exact algorithm in its base version (Rockafellar and Wets, 1991).